Estimate the force you would exert on another person 2 meters away.
Answer #1

Force is given by

F=G\frac{m_1m_2}{r^2}

Force on 1 due to 2 is

F_{12}=G\frac{m_1m_2}{r^2}=G\frac{m_1m_2}{4} N